OpenStudy (anonymous):

How do you find csc(-3pi/4)

OpenStudy (raden):

csc(-3pi/4) = 1/sin(-3pi/4) = 1/(-(sin3pi/4)) = - 1/sin3pi/4 = ... do u know the exact value of sin(3pi/4) ?

OpenStudy (anonymous):

sin(3pi/4) would be \[\frac{ \sqrt{2} }{ 2 }\] right because its the same as sin(pi/4)

OpenStudy (aravindg):

3pi/4=pi-pi/4

OpenStudy (anonymous):

would it just be the negative of the csc(3pi/4) or is that wrong. Please help

OpenStudy (raden):

sin(3pi/4) = sin(pi/4) = 1/2 * sqrt(2)

OpenStudy (anonymous):

@RadEn I got that part my question is how I would make that into csc(-3pi/4)

OpenStudy (raden):

so, what is -1/(1/2 * sqrt(2)) ?

OpenStudy (anonymous):

ok I think I got it now thanks a lot